# Spectroscopy of two-dimensional interacting lattice electrons using symmetry-aware neural backflow transformations

### v1

@@ Abstract
Neural networks have shown to be a powerful tool to represent ground state of quantum many-body systems, including for fermionic systems. In this work, we introduce a framework for embedding lattice symmetries in Neural Slater-Backflow-Jastrow wavefunction ansatzes, and demonstrate how our model allows us to target the ground state and low-lying excited states. To capture the Hamiltonian symmetries, we introduce group-equivariant backflow transformations. We study the low-energy excitation spectrum of the $t$-$V$ model on a square lattice away from half-filling, and find that our symmetry-aware backflow significantly improves the ground-state energies, and yields accurate low-lying excited states for up to $10\times 10$ lattices. We additionally compute the two-point density-correlation function and the structure factor to detect the phase transition and determine the critical point. Finally, we quantify the variational accuracy of our model using the $V$-score.
@@ I Introduction
Strong correlations lead to rich physical phenomena in quantum many-body systems, such as metal-insulator transitions, spin-charge separation, and the paradigmatic fractional quantum hall effect [1, 2, 3, 4]. The strong interactions among particles in these systems make their description complex. Various numerical methods have been developed to tackle the strongly correlated regime, including variational approaches such as variational Monte Carlo (VMC) [5] and tensor network methods [6]. Machine learning has recently found its application in quantum many-body physics to introduce flexible and powerful parameterizations of quantum states. This is guided by the capacity of neural networks to act as universal and efficient high-dimensional function approximators [7]. They have shown great potential, often resulting in state-of-the-art ground state approximations, especially in 2D [8, 9, 10, 11, 12, 13], and have also found their application in dynamics [7, 14, 15, 16, 17, 18, 19, 20].
@@
Neural network quantum states (NQS) [7] have also been used to simulate fermionic systems in the first and second quantization formalisms [21, 22, 23, 24, 25, 26, 27]. In the latter, the fermionic anticommutation relations make variational approaches challenging. This is particularly clear when mapping fermionic operators onto spin operators, e.g. using Jordan-Wigner in >1D, where these mappings introduce a highly non local spin Hamiltonian [28]. On the other hand, in the first quantization formalism one must exactly fulfill the particle-permutation antisymmetry of the wave function. A conventional variational wavefunction typically involves a (mean-field) Slater determinant to account for antisymmetry, combined with a two-body Jastrow factor [29] to capture particle correlations. One way to further improve this ansatz is by introducing correction terms known as backflow transformations (BF). This modification involves making the orbitals within the Slater determinant depend on the positions of all fermions. Feynman and Cohen originally introduced the idea to analyze the excitation spectrum of liquid Helium-4 [30], and was successfully extended to electronic degrees of freedom [31, 32, 33, 34, 35]. Backflow transformations can alter the nodal surface, thereby reducing approximation errors [36, 37, 38, 32]. Recently, the backflow transformation has been introduced as a neural network in the context of NQS applied to discrete [39] and continuous [26, 27] fermionic systems.
For spin degrees of freedom, it has been demonstrated that embedding symmetries into NQS can greatly improve ground state accuracy [40, 21, 41, 42, 13, 43, 44, 45, 46]. Furthermore, restoring the symmetries of the system enables us to target low-lying excited states that can be classified by the different symmetry sectors [40, 21]. One general way to target the low-energy states of the symmetry sectors is by applying quantum-number projectors to the wave function [21].
@@
In this work, we introduce a method for embedding lattice symmetries of 2D fermionic lattice Hamiltonians into neural backflow transformations and demonstrate its efficacy using Slater-Backflow-Jastrow wavefunction ansatzes. We benchmark our ansatz on the $t$-$V$ model on a square lattice and find that it significantly increases the ground-state accuracy compared to other state-of-the-art approaches, and additionally allows us to accurately determine the low-lying excited states over a wide range of interaction strengths.
II Methods
II.1 Fermions on the lattice

### v2

@@ Abstract
Neural networks have shown to be a powerful tool to represent the ground state of quantum many-body systems, including fermionic systems. However, efficiently integrating lattice symmetries into neural representations remains a significant challenge. In this work, we introduce a framework for embedding lattice symmetries in fermionic wavefunctions and demonstrate its ability to target both ground states and low-lying excitations. Using group-equivariant neural backflow transformations, we study the $t$-$V$ model on a square lattice away from half-filling. Our symmetry-aware backflow significantly improves ground-state energies and yields accurate low-energy excitations for lattices up to $10\times 10$. We also compute accurate two-point density-correlation functions and the structure factor to identify phase transitions and critical points. These findings introduce a symmetry-aware framework important for studying quantum materials and phase transitions.
@@ I Introduction
Strong correlations lead to rich physical phenomena in quantum many-body systems, such as metal-insulator transitions, spin-charge separation, and the paradigmatic fractional quantum Hall effect Hubbard_1964, Vijayan_2020, Li_2018, Avella_2013. The strong interactions among particles in these systems make their description complex. Various numerical methods have been developed to tackle the strongly correlated regime, including variational approaches such as variational Monte Carlo (VMC) McMillan_1965 and tensor network methods White_1992. Machine learning has recently found its application in quantum many-body physics to introduce flexible and powerful parameterizations of quantum states. This is guided by the capacity of neural networks to act as universal and efficient high-dimensional function approximators Carleo_2017. They have shown great potential, often resulting in state-of-the-art ground state approximations, especially in 2D Carleo_2019, Deng_2017, Chen_2018, Ferrari_2019, Hibat_2020, Nomura_Imada_2021, and have also found their application in dynamics Carleo_2017, Czischek_2018, Fabiani_2019, Schmitt_2020, Fabiani_2021, Hofmann_2022, nys2024ab, Lin_2022.
@@
Neural network quantum states (NQS) Carleo_2017 have also been used to simulate fermionic systems in the first and second quantization formalisms Choo_2020, Stokes_2020, Hermann_2022, pescia2023message, kim2023neural, Pfau_2020, Hermann_2020. In the latter, the fermionic anticommutation relations make variational approaches challenging. This is particularly clear when mapping fermionic operators onto spin operators, e.g. using Jordan-Wigner in >1D, where these mappings introduce a highly non local spin Hamiltonian Nys_2022. On the other hand, in the first quantization formalism one must exactly fulfill the particle-permutation antisymmetry of the wave function. A conventional variational wavefunction typically involves a (mean-field) Slater determinant to account for antisymmetry, combined with a two-body Jastrow factor Jastrow_1955 to capture particle correlations. One way to further improve this ansatz is by introducing correction terms known as backflow transformations (BF). This modification involves making the orbitals within the Slater determinant depend on the positions of all fermions. Feynman and Cohen originally introduced the idea to analyze the excitation spectrum of liquid Helium-4 Feynman_1956, and it was successfully extended to electronic degrees of freedom holzmann2003backflow, Holzmann_2019, rios2006inhomogeneous, tocchio2008role, tocchio2011backflow. Backflow transformations can alter the nodal surface, thereby reducing approximation errors Kwon_1993, Kwon_1998, Hutcheon_2020, Holzmann_2019. Recently, the backflow transformation has been introduced as a neural network in the context of NQS applied to discrete Luo_2019 and continuous Pfau_2020, Hermann_2020 fermionic systems. For spin degrees of freedom, it has been demonstrated that embedding symmetries into NQS can greatly improve ground state accuracy Lange_2023, Choo_2020, vieijra2021many, vieijra2020restricted, Nomura_Imada_2021, Mizusaka_Imada_2004, Morita_2015, Nomura_2021, Choo_2018. Furthermore, restoring the symmetries of the system enables us to target low-lying excited states that can be classified by the different symmetry sectors Lange_2023, Choo_2020. One general way to target the low-energy states of the symmetry sectors is by applying quantum-number projectors to the wave function Choo_2020.
In this work, we introduce a method for embedding lattice symmetries of 2D fermionic lattice Hamiltonians into neural backflow transformations and demonstrate its efficacy using Slater-Backflow-Jastrow wavefunction ansatzes. Our approach incorporates symmetry-aware neural backflow transformations, fulfilling equivariance conditions for translational and particle-permutation symmetries through convolutional neural networks (CNN). By employing quantum number projection, we symmetrize the wavefunction and accurately target low-lying excited states by varying quantum numbers such as total momentum.
@@
We benchmark our ansatz on the $t$-$V$ model on a square lattice and find that it significantly increases the ground-state accuracy compared to other state-of-the-art approaches. Additionally, it enables precise determination of low-lying excited states over a wide range of interaction strengths. The results demonstrate the robustness of this approach in capturing phase transitions and identifying the critical interaction strength. Furthermore, a V-score analysis highlights the improved variational accuracy of our method across diverse interaction strengths, system sizes, and excitations, showcasing the effectiveness of symmetry-aware neural backflow transformations for fermionic systems.
II Results
II.1 Hamiltonian and Observables
The Hamiltonian of the $t$-$V$ model reads
The first term describes electron hopping between neighboring sites with hopping parameter $t$. The second term corresponds to the nearest neighbor Coulomb repulsion with interaction strength $V\geq 0$. We will set $t=1$ from hereon. We further decompose the Hilbert space $\mathcal{H}$ into fixed particle-number subspaces $\mathcal{H}_{N_{f}}$ Jung_2020. The $t$-$V$ model was originally introduced to study the thermodynamic and transport properties of superconductors Lam_1978, Girvin_1978. Additionally, it provides a conceptual framework for explaining phenomena such as phase separation or stripe order in cuprates and organic conductors Czart_2008, Dagotto_2001, Schulz_1990, Henley_2001, Zhang_2003, Ma_2021. In practice, the $t$-$V$ model can be realized, for example, in experiments with strongly polarized ${}^{3}\mathrm{He}$ atoms  Partridge_2007, Moreo_2007, or using cold atoms in optical lattices with Rydberg dressing Weckesser_2024. Despite its apparent simplicity, the $t$-$V$ model cannot be solved analytically in two or higher dimensions. It also reveals highly nontrivial phase transitions that have been studied in previous works with various computational techniques, including variational Monte Carlo Gubernatis_1985, Scalapino_1984, Czart_2008, de_Woul_2010, Song_2014, Sikora_2015, Corboz_2010, Stokes_2020, Evertz_2003, Dai_2022.
In the strong-coupling limit where $V/t\rightarrow\infty$ we encounter a charge-ordering (CO) phase where the system behaves classically. At large $V/t$, it is energetically unfavorable for two fermions to occupy neighboring lattice sites. The correlations become short-ranged, suggesting localization and insulating behavior, giving rise to a charge-ordered insulating phase. At half-filling, where the particle density is $\bar{n}=0.5$, the charge order forms a checkerboard pattern. Conversely, at weak interaction strengths $V/t$, the fermions can easily hop between neighboring lattice sites, and the system behaves like a non-interacting free Fermi gas. In this weak coupling limit the system enters the metallic phase and becomes exactly solvable at $V/t\rightarrow 0$ Gubernatis_1985, Scalapino_1984, Giamarchi_2010, Stokes_2020, Jordens_2008, Fratini_2009, Greiner_2002, Driscoll_2022, Mocanu_2005.
We introduce the normalized density-density correlation function as Stokes_2020
where $S_{R}(i)=\{j\in\mathcal{V}:d(i,j)=R\}$ is the set of vertices with a fixed distance $R$ from the vertex $i$. Another important observable to detect the CO phase is the structure factor Gubernatis_1985, Scalapino_1984, Song_2014, Sikora_2015, Dang_2008:
In the CO phase, well-defined peaks at $\mathbf{K}=(\pi,\pi)$ indicate a checkerboard charge pattern, and in the thermodynamic limit $S((\pi,\pi))/N_{s}$ converges to a finite value, reflecting long-range order.
We will study the $t$-$V$ model on a two-dimensional square lattice of side $L$ and with periodic boundary conditions, for various system sizes $|\mathcal{V}|=L^{2}$ and different interaction strengths $V/t$, with densities close to half-filling and at closed momentum shells.
II.2 Ground States
We benchmark our symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$) ansatz with respect to a mean-field Slater determinant (which is equivalent to Hartree-Fock (HF)), a symmetrized Slater-Jastrow ($\psi^{K}$) without backflow, and a non-symmetrized Slater-Jastrow. Furthermore, we compare our results with ground state energies obtained using another state-of-the-art neural quantum state method, specifically the "Slater-Jastrow with an additional sign correction neural network" ansatz from Ref. Stokes_2020. In their method, they do not use the symmetry-averaging process. Instead, they employ a Slater-Jastrow inspired ansatz with deep residual networks and convolutional residual blocks to approximately determine the ground state of spinless fermions on a square lattice with nearest-neighbor interaction. The ansatz is a modified Hartree-Fock wavefunction, enhanced by a Jastrow factor and a sign-correcting neural network, both of which are constructed to be invariant under certain lattice symmetries.
It is important to note that the simulations in this section for the different system sizes are performed in an open shell domain, where fermions occupy the lowest-energy available states. In the context of spinless fermions, a “closed shell” configuration occurs when all momentum states within the Fermi surface are fully occupied, typically resulting in a non-degenerate ground state with total momentum $\textbf{K}=(0,0)$. In contrast, an “open shell” configuration arises when some momentum states remain unfilled, leading to a partially filled Fermi surface and potentially different total momenta for the ground state. Thus, in our open shell setup, the ground state does not necessarily correspond to $\textbf{K}=(0,0)$.
In Fig. 1 (a), we present results for a small system size $L=4$ at a density of $\bar{n}=0.44$ and ground state momentum $\textbf{K}=(\pi,0)$, allowing comparison to results obtained by exact diagonalization (ED). Mean-field Slater relative errors range from $10^{-3}$ to $10^{-1}$, with accuracies decreasing at large interaction strengths $V/t$. Our symmetrized backflow ansatz consistently yields ground-state errors below $10^{-3}$ , also at higher values of $V/t$, and achieves lower errors compared to other state-of-the-art methods. We also observe that the symmetry-aware backflow transformation yields the most accurate ground-state energies over the whole interaction regime. For a larger system size $L=10$, at a density of $\bar{n}=0.44$ and ground state momentum $\textbf{K}=(\frac{3\pi}{5},\frac{2\pi}{5})$, shown in Fig. 1 (b), this trend is confirmed. The backflow corrections significantly lower the estimated ground-state energy across all coupling strengths. We compare the converged VMC energies with HF energies by computing the difference $E-E_{\text{HF}}$. Our results show that the backflow ansatz consistently provides lower energies across the full range of interaction strengths.
Figure 1: Comparison of ground-state energies of the $t$-$V$ model obtained with various wavefunction models. We evaluate the performance of the symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$, red) against several other wavefunction models: symmetrized Slater-Jastrow without backflow ($\psi^{K}$, yellow), non-symmetrized mean-field Slater (blue), and non-symmetrized Slater-Jastrow (green). Additionally, we compare to the neural quantum state “Slater-Jastrow with an additional sign correction neural network” described in Ref. Stokes_2020 (black). In panel a), results are presented for a lattice size $L=4$ with $\bar{n}=0.44$, showing relative errors compared to exact diagonalization (ED) as a function of the interaction strengths $V/t$. Panel b) shows the deviation from Hartree-Fock (HF) energies, defined as $E-E_{HF}$, for a system with $L=10$ and a filling fraction $\bar{n}=0.44$. Error bars in both panels represent the corresponding standard deviations, where provided.
In Fig. 2 (a), we show the density-density correlation function as defined in Eq. 2 for $L=10$ close to half-filling $\bar{n}=0.44$ (see Supplementary Note II for results on the $8\times 8$ system). When the interaction strength increases, the correlations start to oscillate more pronouncedly as a result of the increasingly ordered charge distribution. In the CO phase, the amplitudes of the oscillations decrease more gradually with distance compared to that of the metallic phase. For weak couplings, the correlations barely oscillate and fade as the graph distance $R$ increases.
To pinpoint the transition point in the thermodynamic limit, we use finite-size scaling. We study the structure factor $S(\pi,\pi)/N_{s}$ for various $V/t$. The critical transition point is found where the structure factor attains a finite value in the thermodynamic limit. In Fig. 2 (b) we plot the structure factor $S(\pi,\pi)/N_{s}$ as a function of the inverse system size $1/L$ (for $L=4,6,8,10$) and in Table 1 we report the extrapolated results in the thermodynamic limit.
We apply a least squares regression using the SciPy Scipy_2020 library to fit the data and extrapolate $S(\pi,\pi)/N_{s}$ to the thermodynamic limit. In the metallic phase, the structure factor is expected to vanish for large system sizes, such that there exists a finite critical system size, $L_{c}$, beyond which the order parameter vanishes. This reflects the short-range nature of correlations in the metallic phase, which become negligible as the system size increases, causing the structure factor to diminish and ultimately disappear. In contrast, in the CO phase, $L_{c}$ tends to infinity, meaning the structure factor remains finite in the thermodynamic limit. The critical interaction strength, $V_{c}/t$, was identified at the point where the extrapolated $S_{\infty}$ transitions from the metallic phase to a positive finite value in the CO phase, signaling the onset of long-range order. Error estimates for $V_{c}/t$ were calculated using error propagation, based on the uncertainties in $S_{\infty}$ at these transition points.
Near half-filling, specifically for $\bar{n}=0.44$, we estimate that the transition occurs at $V_{c}/t\simeq 1.14\pm 0.04$, which is consistent with the value reported in Ref. Stokes_2020. While the order parameter from Eq. (3) of Ref. Stokes_2020 explicitly assumes a symmetry-broken ground state, we enforce the symmetry and nevertheless find a transition point compatible with Ref. Stokes_2020 using the density-density structure factor. We selected a density close to half-filling to observe the metallic to CO phase transition while avoiding the need for interpolating densities. This choice allowed us to explore conditions slightly doped away from half-filling, given that the phases at exact half-filling have already been extensively studied Gubernatis_1985, Scalapino_1984, Czart_2008, Song_2014, Sikora_2015, Stokes_2020.
Figure 2: Two-point correlations and finite-size scaling of the structure factor. a) The two-point density-density correlation functions as a function of graph distance $R$ for different interaction strengths $V/t$ for a system size of $L=10$ and particle density $\bar{n}=0.44$. b) Finite-size scaling of the structure factor $S(\pi,\pi)/N_{s}$ versus different system sizes $1/L$ (where $L=4,6,8,10$) with $\bar{n}=0.44$ for different values of $V/t$ close to the estimated critical point $V_{c}/t\simeq 1.14\pm 0.04$. Error bars represent the corresponding standard deviations.
$V/t$ $S_{\infty}$
0.1 $0.0$
0.6 $0.0$
1.0 $0.0$
1.3 $0.003\pm 0.001$
1.6 $0.042\pm 0.002$
2.7 $0.085\pm 0.002$
Table 1: Thermodynamic limit of the structure factor. Structure factor extrapolations to the thermodynamic limit (taken from Figure 2 (b)) as a function of interaction $V/t$. Values in the metallic phase have been adjusted to reflect physical zeros for consistency.
II.3 Excitations
[…]
Figure 3: High-symmetry points in the Brillouin zone. Single quadrant of the first Brillouin zone of the two-dimensional square lattice, highlighting the high-symmetry points $(k_{u},k_{v})$.
To capture low-lying excitations, we carry out VMC optimizations across various momentum sectors $\mathbf{K}=(k_{u},k_{v})$, where $k_{u,v}\in\left\{0,\pm\frac{2\pi m}{L_{u,v}}\,\middle|\,m=1,2,\dots,\frac{L_{u,v}}{2}\right\}$. Here $L_{u}$ and $L_{v}$ are the side lengths of the two-dimensional lattice in the $\mathbf{e}_{u}$ and $\mathbf{e}_{v}$ directions. In Fig. 3, we represent a single quadrant within the corresponding first Brillouin zone with conventional symbols to represent high-symmetry points  Khanikaev_2005.
Figure 4: Energy comparison and relative errors for closed-shell $L=4$. a) The lowest ($E_{0}^{K}$, teal horizontal lines) and second-lowest energies ($E_{1}^{K}$, yellow horizontal lines) for the designated $\mathbf{K}$ sector using exact diagonalization (ED), compared to the lowest energies in each sector obtained with our variational ansatzes: symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$, blue) and symmetrized Slater-Jastrow without backflow ($\psi^{K}$, red). b) The relative energy errors associated with the ansatzes with respect to the lowest energy in each sector. These results are for a closed-shell system with lattice size $L=4$ and a density of $\bar{n}=0.31$. Error bars represent the corresponding standard deviations.
We first benchmark the performance of our symmetrized Neural Slater-Backflow-Jastrow variational ansatz ($\psi^{K}_{BF}$) and a symmetrized Slater-Jastrow ($\psi^{K}$) (top) model for a system size of $L=4$ and $\bar{n}=0.31$ in Fig. 4. We observe that also in symmetry sectors different from the ground state, the symmetric backflow transformation reduces the relative energy error. We show the lowest energy state in each sector and compare it with results obtained from ED. In the lower panel of Fig. 4, we calculate the corresponding relative errors with respect to the ED energies. For our backflowed ansatz ($\psi^{K}_{BF}$) these lie between $10^{-6}-10^{-3}$ for all sectors. In Supplementary Note IV, we also include additional simulation data of the open shell $L=4$ system (see Fig. S3).
Prior studies have documented a range of coexistence phenomena for large and finite $V/t$ away from half-filling, transitioning from phase separation to potential stripe and checkerboard coexistence. In Ref. Chen_2010, the $t$-$V$ model on a square lattice with nearest-neighbor repulsive interactions was studied using mean-field theory for small system sizes. The authors observed a second-order phase transition from the Fermi liquid to the $(\pi,\pi)$ charge density wave state. At stronger repulsion, charge density waves coexisted at different momentum sectors when doped away from half-filling. In Ref. Henley_2001 and the subsequent study in Ref. Zhang_2003, exact diagonalization was used to study small 2D systems. They found that at high repulsion and around quarter-filling densities, doped holes formed stable charged stripes acting as anti-phase walls mila1994exact, which are stable against phase separation in fermionic systems.
We extend our analysis to larger system sizes $L=8$ with $\bar{n}=0.39$ and $L=10$ with $\bar{n}=0.41$ to reduce finite-size effects. We simulate each system size with distinct particle densities to ensure they remain within the closed-shell domain, avoiding ground-state degeneracies in the non-interacting limit. We confirm the persistent non-degenerate ground state at the $\Gamma=(0,0)$ point (see Fig. 5 for $L=8$ and Fig. S2 in Supplementary Note III for $L=10$). By comparing our symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$) ansatz with a symmetrized Slater-Jastrow ($\psi^{K}$) without backflow, we observe improved energy levels, even for low-lying excited states, with the inclusion of the backflow correction term.
Figure 5: Lowest excitation energies for $L=8$ closed-shell system. Lowest excitation energies in different $\mathbf{K}$ sectors for a lattice of size $L=8$, corresponding to a closed-shell configuration with particle density $\bar{n}=0.39$, evaluated across various interaction strengths $V/t$. Results are shown for symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$, blue) and symmetrized Slater-Jastrow without backflow ($\psi^{K}$, red). Error bars represent the corresponding standard deviations.
[…]
II.4 V-score
We now aim to generalize the assessment of the performance of our model. Since ED becomes intractable for larger system sizes, we rely on the recently introduced V-score as a guiding metric Wu_2023, which can be computed using the variational energy and its variance. The V-score is dimensionless and invariant under energy shifts by construction. It is defined as Wu_2023:
The V-score serves as a valuable tool for discerning which Hamiltonians and regimes pose challenges for arbitrary classical variational techniques, even when we lack prior knowledge about the precise exact solution. Its practicality lies in its ability to quantify the accuracy of a particular method independently, without the need for direct comparisons with other methods. The intuition behind the V-score is that the energy variance, which becomes zero for an exact ground state, serves as a direct measure of how close a variational state is to the true ground state. This allows us to infer the accuracy of a method based on the variance alone. In particular, this metric enables us to draw comparisons between the accuracy obtained with our method on the given Hamiltonian, compared to other commonly studied condensed matter Hamiltonians, including spin Hamiltonians.
Figure 7: V-score comparison for different variational ansatzes. Ground state V-scores as a function of interaction strength $V/t$ are compared for different ansatzes: $\psi^{K}_{BF}$ (symmetrized backflow Slater-Jastrow, blue/green), $\psi^{K}$ (symmetrized Slater-Jastrow, violet/orange), and HF (Hartree-Fock, red/yellow), shown for system sizes $L=4$ with particle density $\bar{n}=0.31$ (triangles) and $L=10$ with $\bar{n}=0.41$ (stars).
Figure 8: V-score comparison across system sizes and states. V-score for a symmetrized Neural Slater-Backflow-Jastrow as a function of interaction strengths $V/t$ for ground (GS, triangles) and excited states (stars). Data are presented for different system sizes $L$ and particle densities $\bar{n}$: $L=4$ with $\bar{n}=0.31$ (blue), $L=8$ with $\bar{n}=0.39$ (yellow), and $L=10$ with $\bar{n}=0.41$ (red). Note that excitation data are available for $V/t$ values in the set {0.0, 0.1, 1.0, 10.0}, while ground-state data cover a broader range, including {0.0, 0.1, 1.0, 2.0, 4.0, 6.0, 10.0}.
III Discussion
In this work, we introduced an approach to studying the low-energy excitation spectrum of fermionic Hamiltonians. By introducing symmetry-aware neural backflow transformations, we show that we can target the eigenstates of fermionic Hamiltonians with high accuracy. As a benchmark comparison, we show that this approach also yields significantly more accurate ground-state energies than other state-of-the-art variational Monte Carlo approaches.
In particular, we introduce equivariance conditions for the backflow transformations that lead to an efficient symmetry projection. We show that convolutional neural networks yield a powerful parametrization for our symmetry-aware backflow transformations that fulfill the equivariance conditions for both translation and particle-permutation symmetry. This key contribution enables us to efficiently access excited states by varying the total momentum $\mathbf{K}$ in the quantum number projection. Furthermore, we have showcased the utility of our approach in identifying phase-transitions on the $t$-$V$ model at system sizes far beyond what is reachable with exact diagonalization. To this end, we computed correlation functions and structure factors, resulting in the pinpointing of the critical point at $V_{c}/t=1.14$. We also computed the V-score to quantify the variational accuracy of our proposed ansatz for different interaction regimes, system sizes, and excitations. Previous analysis based on the V-score has highlighted the challenging nature of targeting fermionic eigenstates. Our observations indicate that the symmetry-aware backflow ansatz yields accurate ground states and performs favorably compared to other state-of-the-art methods over the full interaction regime, including when strong correlations occur. Additionally, we find that our method yields accurate approximations to the low-energy eigenstates with a given $K$ momentum.
Future extensions of our approach include generalizing it by including additional symmetries, such as rotational and reflection symmetries. This will involve using more general group-convolutional kernels, as in group-convolutional neural networks (GCNN) Cohen_2016, Roth_2021, Roth_2023. This would require addressing the increased computational demands as the number of symmetry elements grows. We focused on the nearest neighbor $t$-$V$ model, but an extension is to consider Hamiltonians where spin-degrees of freedom become relevant as well (such as the Fermi-Hubbard model), or where interactions beyond nearest neighbor terms become relevant.
IV Methods
IV.1 Fermions on the lattice